# A two-phase model of galaxy formation: III. The formation of globular clusters

### v2

@@ Abstract
We develop a model of globular cluster (GC) formation within the cosmological hierarchy of structure formation. The model is rooted in the ‘two-phase’ scenario of galaxy formation developed in Paper-I, where the fast accretion of dark matter halos at high redshift leads to the formation of self-gravitating, turbulent gas clouds that subsequently fragment into dynamically hot systems of dense sub-clouds with masses $\sim 10^{6}$$10^{7}{\rm M_{\odot}}$. Here we elaborate on the formation, evolution, and fate of these sub-clouds, and show that some of the sub-clouds can be compactified via two distinctive channels into a ‘supernova-free’ regime to form two distinct populations of GCs. The model is simple, characterized by a small number of free parameters underpinned by physical considerations, and can be efficiently implemented into cosmological N-body simulations to generate a coherent sample of halos, galaxies, and GCs. Our model can reproduce a range of observations on GCs, including the mass function, the size-mass relation, the frequency per unit host galaxy/halo mass, the bimodal metallicity distribution, and the spatial profile. Predictions for GCs are made for both the local Universe and for redshift up to $z\approx 10$, and can be tested by upcoming observations.
@@ 1 Introduction
The main challenge in modeling GC formation in a cosmological context is the large gap between the properties of GCs and their host galaxies. GCs are relatively small, with a typical size of $\sim 10\,{\rm pc}$ (e.g. Krumholz et al., 2019) that is about two orders of magnitude smaller than a typical dwarf galaxy (e.g. Wang et al., 2023). Consequently, certain approximations are necessary to simulate GC formation as a part of galaxy formation in the cosmic density field. Along this line, two semi-analytic approaches have been developed, each incorporating different physical inputs derived from numerical simulations. The first approach involves post-processing cosmological hydrodynamic simulations, or their zoom-in runs, with subgrid physics specific to GC formation. Depending on the resolution of the simulation used, the input conditions for GC formation can range from global properties of individual GMCs (e.g. Grudić et al., 2023) to the smoothed field of the gas properties (e.g. Kruijssen et al., 2011; Pfeffer et al., 2018; Kruijssen et al., 2019a). The second approach more closely resembles traditional semi-analytic models of galaxies based on dark matter halos and their merger trees. The global properties of individual galaxies are obtained from a semi-analytic model (El-Badry et al., 2019; De Lucia et al., 2023) or an empirical model of galaxy formation (Li & Gnedin, 2014; Choksi et al., 2018; Valenzuela et al., 2021), while the criteria for GC formation are based on some properties of the assembly histories of galaxies/halos, such as mergers of galaxies (El-Badry et al., 2019) and rates of halo mass assembly (Li & Gnedin, 2014; Choksi et al., 2018; Valenzuela et al., 2021; Chen & Gnedin, 2022). Both approaches have been successful in reproducing some observed properties of GCs. Despite the progress, significant knowledge gaps still exist in our understanding of the GC population. Indeed, our understanding of galaxy formation, in particular of star formation and feedback processes, is still incomplete, and so the large-scale conditions for the formation of GCs are still uncertain. In addition, small-scale processes driving the formation and disruption of GCs are also not well understood.
@@
Recent observations by the James Webb Space Telescope (JWST) have opened a new avenue to study galaxy formation in the early Universe. Recent observations indicated that early galaxies may differ significantly from their counterparts in the local Universe, suggesting that the processes driving galaxy formation at high $z$ may be different from those in the low-$z$ Universe. In particular, the efficiency of gas assumption by star formation seems to be enhanced significantly at high $z$ (Bouwens et al., 2023; Finkelstein et al., 2023; Xiao et al., 2023; Weibel et al., 2024; Wang et al., 2024), supermassive black holes (SMBHs) appear to be over-massive for their host galaxy mass (Maiolino et al., 2023; Pacucci et al., 2023; Matthee et al., 2024; Pacucci & Loeb, 2024), there seems to be an inversion of metallicity gradients in both the ISM and circumgalactic medium (CGM; Wang et al., 2022; Lin et al., 2023), and the total mass of bound star clusters in galaxies seems to be excessively large (Vanzella et al., 2022, 2023; Welch et al., 2023; Lin et al., 2023; Claeyssens et al., 2023; Adamo et al., 2024; Messa et al., 2024; Fujimoto et al., 2024; Mowla et al., 2024; Whitaker et al., 2025). Given that many GCs are old, these discoveries at high $z$ have important implications for understanding GC formation.
@@
This paper is organized as follows. In §2, we describe the context of galaxy formation within which GCs emerge. In §3, we elaborate on our model of GC formation. In §4, we describe numerical implementations of our model, present its predictions and compare them with observations and with other models. In §5, we summarize our results and discuss their implications. Throughout this paper, we adopt a flat $\Lambda$CDM cosmology with parameters derived from the Planck2015 results (Planck Collaboration et al., 2016): the density parameters $\Omega_{\rm M,0}=0.3089$, $\Omega_{\rm B,0}=0.0486$, and $\Omega_{\rm\Lambda,0}=0.6911$; the Hubble constant $H_{0}=100\,h\,{\rm km\,s^{-1}\,Mpc^{-1}}$, with $h=0.6774$; the Gaussian initial density field with a power spectrum $P(k)\propto k^{n}$ with $n=0.9667$; and an amplitude specified by $\sigma_{8}=0.8159$. To avoid confusion, we use ‘log’ to denote the base-10 logarithm, and ‘ln’ to denote the natural logarithm. We use the terms 1, 2, and 3-$\sigma$ ranges to denote those centralized at the median value and covering $68\%$, $95\%$, and $99.7\%$, respectively, of the probability mass for any probability distribution.
@@ 2 The context of galaxy formation
Our goal is to develop a paradigm to understand the formation of GCs in the current cosmological framework of galaxy formation. Due to the large dynamic range covered by our model, we structure it in a hierarchical way, from large scales of low density to small scales of high density. The basic idea behind the hierarchy is to introduce different physical and numerical models to describe the ‘jumps’ in characteristic scales, and join them with recipes describing their interfaces. Thus, we start with modeling the structure formation at large scales, and use the results as boundary conditions, i.e. ‘environments’, for the next step on smaller scales. The first scale, describing the large-scale structures of the cosmic density field and its interface with the second scale, which describes dark matter halos, is modeled using results from cosmological N-body simulations. Models of galaxy formation and the growth of SMBH are built on halos and their merger trees, using the two-phase paradigm described in Paper-I, which bridges the interface between halos and galaxies. Gas fragmentation and star formation are modeled by a new recipe developed in this paper, which bridges the gap to the fourth scale, sub-clouds in galaxies, where globular clusters form. Fig. 1 shows the overall context of our model, where the two-point auto-correlation function of globular clusters, predicted by our model, is shown to represent the scales and densities of the structures in question. The spatial and density scales of relevant gas processes (e.g. mass accretion, gas cooling, fragmentation and compression) and the corresponding structures (large-scale structure, halos, galaxies, and sub-clouds) are marked for reference. This plot demonstrates schematically how the formation of GCs fits in the general cosmological context, and how the related physical processes define the characteristic scales involved in our problem. In what follows, we describe these processes in more detail.
@@ 3 The formation of globular clusters
A main conclusion of our two-phase model (Paper-I) is the formation of sub-structures, referred to as sub-clouds, once the SGC becomes self-gravitating. In the fast phase, the SGC is expected to be gas-rich and turbulent due to the fast accretion of halo. High-speed winds from SN and AGN can also contribute to the turbulence (Ma et al., 2020; Mercedes-Feliz et al., 2023). Significant compression of the sub-clouds is expected if the turbulence is supersonic, which may produce dense sub-clouds for the formation of GCs we are interested in here. Despite the complexity of the details, it is generally believed that the properties of sub-clouds in a turbulent ISM follow ‘quasi-universal’ laws and distributions, as suggested by both observations (Larson, 1981; Lada & Dame, 2020) and numerical simulations (Burkhart et al., 2017; Ma et al., 2020; Grudić et al., 2021; Appel et al., 2022; Buck et al., 2022; Kiihne et al., 2023). In this section, we sample the sub-cloud population using these universal distributions under the constraints given by the global properties of SGCs. We then derive the condition for star formation in each sub-clouds, taking into account both external and internal feedback effects on the sub-clouds. We demonstrate that the condition for a sub-cloud to form a GC leads to two distinct channels for GC formation. Finally, we model the disruption and survival of GCs, which is needed to make predictions for comparisons with observations.
@@ 3.1 The formation of sub-clouds
    Positive feedback. High-speed winds driven by feedback processes can also drive turbulence and compress sub-clouds to high density. Numerical simulations show that quasar winds (Mercedes-Feliz et al., 2023) and SN winds (Ma et al., 2020) can both lead to the formation of dense stellar clumps reminiscent of YMSCs and GCs. The recent observations by JWST (Pereira-Santaella et al., 2024) and ALAM (Roman-Oliveira et al., 2024) also support the presence of stellar feedback as the driver of turbulence. This channel of turbulence formation exists as long as star formation and/or AGN is active, and thus may cover the entire history of a galaxy.
@@
where $s=\ln(\rho_{\rm sc}/\rho_{0})$ is the logarithmic density contrast, $\rho_{0}$ is the mean density of the SGC, $\sigma_{\rm s}$ is a free parameter defining the width of the log-normal distribution, $s_{0}=(-1/2)\sigma_{\rm s}^{2}$ is the median, and $\alpha_{\rm s}$ is the power-law index of the high-density tail. The log-normal piece is found to well represent the isothermal density distribution of a low-density, turbulent ISM (e.g. Burkhart et al., 2009; Myers, 2015). The power-law tail at the high density, on the other hand, is expected to originate from compression by self-gravitating collapses and feedback (e.g. Slyz et al., 2005; Collins et al., 2011; Burkhart et al., 2017; Khullar et al., 2021; Appel et al., 2022; Buck et al., 2022; Appel et al., 2022; Kiihne et al., 2023). The continuity of $p(s)$ and its first derivative at $s=s_{\rm t}$ gives
@@
    $\alpha_{\rm s}$, the high-density slope, is found to be 11112222 by numerical simulations (e.g. Appel et al., 2022; Kiihne et al., 2023). We take $\alpha_{\rm s}=1.5$, which is consistent with the analytical expectation for the pressure-free (free-fall) collapse (Girichidis et al., 2014), and for the weak magnetic field and high compression we assumed in setting $f_{\rm s}$ (Collins et al., 2011).
@@ 3.2 The cooling and fragmentation of sub-clouds
Grey shades in Fig. 6 show the distributions of sub-cloud density at formation ($n_{\rm sc}$), for all sub-clouds formed in the main branches of central subhalos at $z=0$ with different halo masses. Depending on internal properties and environments, sub-clouds can have different fates: some will be destroyed by external feedback, while the others will continue to contract, fragment and form stars at $n_{\rm sc}^{\prime}$. Only the most massive and/or densest sub-clouds can shield themselves from feedback and form sufficient amounts of stars to qualify as (the progenitors of) globular clusters. These will be discussed thoroughly in §§3.3, 3.4 and 3.5, and be synthesized into a coherent model in §4.1. Briefly, the aforementioned conditions, together with supersonic turbulence and metal-poor environment, give rise to two distinct formation channels that lead to two populations of GCs: Pop-I (the metal-rich population) and Pop-II (the metal-poor population).
@@ 3.4 Star formation in sub-clouds
This efficiency reaches unity at $\Sigma_{\rm sc}\approx 10^{3.2}\,{\rm M_{\odot}}{\rm pc}^{-2}$, comparable to the condition for sub-clouds with $\epsilon_{\rm w}\approx 1$. At high metallicity, the momentum injection from multiple scattering can effectively reduce the star formation. A more general form of Eq. (55) was originally proposed by Fall et al. (2010), and then confirmed in subsequent hydrodynamical simulations and observations. In particular, Grudić et al. (2018) showed that the formula is accurate for a mixture of feedback processes, both momentum-driven and energy-driven, including radiation pressure, photo-ionization, stellar winds, etc. The formula is also consistent with various observations of proto-clusters/sub-clouds in the MW and the Large Magellanic Cloud (LMC; Mok et al., 2021).
@@ 4.2 Mass and age distributions
Figure 9: Mass functions of GCs. Different panels show the mass functions of GCs in central galaxies with different halo masses at $z=0$. Both in-situ and ex-situ GCs are included. In each panel, grey curve shows the birth-time mass function of all GCs formed in the history, while black curve shows the $z=0$ mass function of GCs survived after the mass loss due to dynamical evolution is incorporated. Red and blue curves show the Pop-I and Pop-II sub-components, respectively, of the black curve. See §4.1 for the details of modeling and sample selection. In the center panel, the histograms show the observed GC mass functions for MW (green) and M31 (purple), compiled by Rodriguez et al. (2023). Orange markers show the mass function of star clusters in M51, obtained from the star-cluster catalog of LEGUS (Messa et al., 2018). See §4.2 for a detailed discussion of this figure.
Figure 10: Age distributions of GCs. Different panels show the age distributions of GCs in central galaxies with different halo masses at $z=0$. Both in-situ and ex-situ GCs are included. In each panel, grey curve shows the distribution of the birth time of all GCs formed in the history, while black curve shows that of GCs survived until $z=0$ after the mass loss due to dynamical evolution is incorporated. Red and blue curves show the Pop-I and Pop-II sub-components, respectively, of the black curve. See §4.1 for the details of modeling and sample selection. In the center panel, the histograms show the results of the empirical model of Chen & Gnedin (2024) for MW (green) and M31 (purple). Their results for all and survived GCs are shown by dashed and solid histograms, respectively. Orange markers show the age distribution of star clusters more massive than $10^{5}{\rm M_{\odot}}$ in M51, obtained from the star-cluster catalog of LEGUS (Messa et al., 2018). See §4.2 for a detailed discussion of this figure.
@@
In the center panel of Fig. 9, we also include the observed mass functions of GCs in the Milky Way (green curve) and M31 (purple curve), compiled by Rodriguez et al. (2023, see their figure 3) using catalogs from Harris (1996, 2010) and Caldwell et al. (2011), respectively. The observed GC mass function of M31 has an amplitude about $0.5\,{\rm dex}$ higher than that of the MW, and appears more top-heavy. This may be explained partly by the more massive halo of M31, and partly by the fact that it has more recent mergers that can bring in ex-situ GCs as well as increase the in-situ formation of GCs (see the age distribution below; see also, e.g. §4.3 of Chen & Gnedin, 2024). The mass function of GCs predicted by our model (black curve) falls between those of MW and M31 over the entire range of $M_{\rm cls}$, indicating that the model is capable of reproducing the average GC mass function for MW-size galaxies.
@@
The difference between MW and M31 in the mass and age distributions of GCs indicates the diversity of GC formation in galaxies with different states and assembly histories. To clearly see this, we consider another galaxy, M51 (NGC 5194), a spiral galaxy interacting with a massive companion, NGC 5195. The stellar mass of M51 is inferred to be $\approx 10^{10.6}{\rm M_{\odot}}$ (Leroy et al., 2008; Wei et al., 2021), slightly smaller than that of MW. We take the star-cluster candidates of the LEGUS survey for M51 (Messa et al., 2018), remove the contamination by selecting only those with machine-learning class 1 or 2, and exclude those with age less than $10\,{\rm Myr}$ to avoid unbound systems. The mass distribution of the star clusters so obtained is shown in the center panel of Fig. 9. We further restrict the sample to massive clusters with $M_{\rm cls}\geqslant 10^{5}{\rm M_{\odot}}$ so that they are age-complete (see Fig. 5 of Messa et al., 2018), and show the age distribution of these clusters in the center panel of Fig. 10. At $M_{\rm cls}\gtrsim 10^{5}{\rm M_{\odot}}$, the cluster mass function of M51 is much lower than those of MW, M31 and the model prediction. The age distribution shows that these massive clusters are mostly born at $z<0.5$ by recent star formation, and none of them are as old as those produced at $z\approx 5$, the peak of the GC-formation activity predicted by our model. Such a situation can be created if M51 forms very late, and is hosted by a halo with very stable dynamical environment in the early stage. The lack of galactic-scale turbulence thus suppresses the formation of old and massive GCs, and large amounts of GCs are produced only later by gas compression through, e.g. disk instabilities and/or external interactions (see the discussion in §3.1).

### v3

@@ Abstract
We develop a model of globular cluster (GC) formation within the cosmological hierarchy of structure formation. The model is rooted in the ‘two-phase’ scenario of galaxy formation developed in Paper-I, where the fast accretion of dark matter halos at high redshift leads to the formation of self-gravitating, turbulent gas clouds that subsequently fragment into dynamically hot systems of dense sub-clouds with masses $\sim 10^{6}$$10^{7}{\rm M_{\odot}}$. Here we elaborate on the formation, evolution, and fate of these sub-clouds, and show that some of the sub-clouds can be compactified via two distinctive channels into a ‘supernova-free’ regime to form two distinct populations of GCs. The model is simple, characterized by a small number of free parameters underpinned by physical considerations, and can be efficiently implemented into cosmological N-body simulations to generate a coherent sample of halos, galaxies, and GCs. Calibrated with observations, our model can reproduce a range of observational statistics, including those for GC masses, sizes, metallicities, spatial distributions, and the relation of GC systems with host galaxies/halos. Significant discrepancies between model results and existing observations are discussed in connection to processes implemented in the model. Predictions for GCs are made for both the local Universe and for redshift up to $z\approx 10$, and can be tested by upcoming observations.
@@ 1 Introduction
The main challenge in modeling GC formation in a cosmological context is the large gap between the properties of GCs and their host galaxies. GCs are relatively small, with a typical size of $\sim 10\,{\rm pc}$ (e.g. Krumholz et al., 2019) that is about two orders of magnitude smaller than a typical dwarf galaxy (e.g. Wang et al., 2023). Consequently, certain approximations are necessary to simulate GC formation as a part of galaxy formation in the cosmic density field. Along this line, two semi-analytic approaches have been developed, each incorporating different physical inputs derived from numerical simulations. The first approach involves post-processing cosmological hydrodynamic simulations, or their zoom-in runs, with subgrid physics specific to GC formation. Depending on the resolution of the simulation used, the input conditions for GC formation can range from global properties of individual GMCs (e.g. Grudić et al., 2023) to the smoothed field of the gas properties (e.g. Kruijssen et al., 2011; Pfeffer et al., 2018; Kruijssen et al., 2019a). The second approach more closely resembles traditional semi-analytic models of galaxies based on dark matter halos and their merger trees. The global properties of individual galaxies are obtained from a semi-analytic model (El-Badry et al., 2019; De Lucia et al., 2024) or an empirical model of galaxy formation (Li & Gnedin, 2014; Choksi et al., 2018; Valenzuela et al., 2021), while the criteria for GC formation are based on some properties of the assembly histories of galaxies/halos, such as mergers of galaxies (El-Badry et al., 2019) and rates of halo mass assembly (Li & Gnedin, 2014; Choksi et al., 2018; Valenzuela et al., 2021; Chen & Gnedin, 2022). Both approaches have been successful in reproducing some observed properties of GCs. Despite the progress, significant knowledge gaps still exist in our understanding of the GC population. Indeed, our understanding of galaxy formation, in particular of star formation and feedback processes, is still incomplete, and so the large-scale conditions for the formation of GCs are still uncertain. In addition, small-scale processes driving the formation and disruption of GCs are also not well understood.
@@
Recent observations by the James Webb Space Telescope (JWST) have opened a new avenue to study galaxy formation in the early Universe. These observations indicated that early galaxies may differ significantly from their counterparts in the local Universe, suggesting that the processes driving galaxy formation at high $z$ may be different from those in the low-$z$ Universe. In particular, the efficiency of gas consumption by star formation seems to be enhanced significantly at high $z$ (Bouwens et al., 2023; Finkelstein et al., 2023; Xiao et al., 2024; Weibel et al., 2024; Wang et al., 2024); supermassive black holes (SMBHs) appear to be over-massive for their host galaxy mass (Pacucci et al., 2023; Maiolino et al., 2024; Matthee et al., 2024; Pacucci & Loeb, 2024); there seems to be an inversion of metallicity gradients in both the ISM and circumgalactic medium (CGM; Wang et al., 2022; Lin et al., 2023); and the total mass of bound star clusters in galaxies seems to be excessively large (Vanzella et al., 2022, 2023; Welch et al., 2023; Lin et al., 2023; Claeyssens et al., 2023; Adamo et al., 2024; Messa et al., 2024; Fujimoto et al., 2024; Mowla et al., 2024; Whitaker et al., 2025). Given that many GCs are old, these discoveries at high $z$ have important implications for understanding GC formation.
@@
This paper is organized as follows. In §2, we describe the context of galaxy formation within which GCs emerge. In §3, we elaborate on our model of GC formation. In §4, we describe numerical implementations of our model, present its predictions and compare them with observations and with other models. In §5, we summarize our results and discuss their implications. Throughout this paper, we adopt a flat $\Lambda$CDM cosmology with parameters derived from the Planck2015 results (Planck Collaboration et al., 2016): the density parameters $\Omega_{\rm M,0}=0.3089$, $\Omega_{\rm B,0}=0.0486$, and $\Omega_{\rm\Lambda,0}=0.6911$; the Hubble constant $H_{0}=100\,h\,{\rm km\,s^{-1}\,Mpc^{-1}}$, with $h=0.6774$; the Gaussian initial density field with a power spectrum $P(k)\propto k^{n}$ with $n=0.9667$; and a perturbation amplitude specified by $\sigma_{8}=0.8159$. To avoid confusion, we use ‘log’ to denote the base-10 logarithm, and ‘ln’ to denote the natural logarithm. We use the terms 1, 2, and 3-$\sigma$ ranges to denote those centralized at the median value and covering $68\%$, $95\%$, and $99.7\%$, respectively, of the probability mass for any probability distribution.
@@ 2 The context of galaxy formation
Our goal is to develop a paradigm to understand the formation of GCs in the current cosmological framework of galaxy formation. Due to the large dynamic range covered by our model, we structure it in a hierarchical way, from large scales of low density to small scales of high density. The basic idea behind the hierarchy is to introduce different physical and numerical models to describe the ‘jumps’ in characteristic scales, and to join them with recipes describing their interfaces. Thus, we start with modeling the structure formation at large scales, and use the results as boundary conditions, i.e. ‘environments’, for the next step on smaller scales. The first scale, describing the large-scale structures of the cosmic density field and its interface with the second scale, which describes dark matter halos, is modeled using results from cosmological N-body simulations. Models of galaxy formation and the growth of SMBH are built on halos and their merger trees, using the two-phase paradigm described in Paper-I, which bridges the interface between halos and galaxies. Gas fragmentation and star formation are modeled by a new recipe developed in this paper, which bridges the gap to the fourth scale, sub-clouds in galaxies, where globular clusters form. Fig. 1 shows the overall context of our model, where the two-point auto-correlation function of globular clusters, predicted by our model, is shown to represent the scales and densities of the structures in question. The spatial and density scales of relevant gas processes (e.g. mass accretion, gas cooling, fragmentation and compression) and the corresponding structures (large-scale structure, halos, galaxies, and sub-clouds) are marked for reference. This plot demonstrates schematically how the formation of GCs fits in the general cosmological context, and how the related physical processes define the characteristic scales involved in our problem. In what follows, we describe these processes in more detail.
@@ 3 The formation of globular clusters
A main conclusion of our two-phase model (Paper-I) is the formation of sub-structures, referred to as sub-clouds, once the SGC becomes self-gravitating. In the fast phase, the SGC is expected to be gas-rich and turbulent due to the fast accretion of halo. High-speed winds from SN and AGN can also contribute to the turbulence (Ma et al., 2020; Mercedes-Feliz et al., 2024). Significant compression of the sub-clouds is expected if the turbulence is supersonic, which may produce dense sub-clouds for the formation of GCs we are interested in here. Despite the complexity of the details, it is generally believed that the properties of sub-clouds in a turbulent ISM follow ‘quasi-universal’ laws and distributions, as suggested by both observations (Larson, 1981; Lada & Dame, 2020) and numerical simulations (Burkhart et al., 2017; Ma et al., 2020; Grudić et al., 2021; Appel et al., 2022; Buck et al., 2022; Kiihne et al., 2025). In this section, we sample the sub-cloud population using these universal distributions under the constraints given by the global properties of SGCs. We then derive the condition for star formation in each sub-clouds, taking into account both external and internal feedback effects on the sub-clouds. We demonstrate that the condition for a sub-cloud to form a GC leads to two distinct channels for GC formation. Finally, we model the disruption and survival of GCs, which is needed to make predictions for comparisons with observations.
@@ 3.1 The formation of sub-clouds
    Positive feedback. High-speed winds driven by feedback processes can also drive turbulence and compress sub-clouds to high density. Numerical simulations show that quasar winds (Mercedes-Feliz et al., 2024) and SN winds (Ma et al., 2020) can both lead to the formation of dense stellar clumps reminiscent of YMSCs and GCs. The recent observations by JWST (Pereira-Santaella et al., 2024) and ALMA (Roman-Oliveira et al., 2024) also support the presence of stellar feedback as the driver of turbulence. This channel of turbulence formation exists as long as star formation and/or AGN is active, and thus may cover the entire history of a galaxy.
@@
where $s=\ln(\rho_{\rm sc}/\rho_{0})$ is the logarithmic density contrast, $\rho_{0}$ is the mean density of the SGC, $\sigma_{\rm s}$ is a free parameter defining the width of the log-normal distribution, $s_{0}=(-1/2)\sigma_{\rm s}^{2}$ is the median, and $\alpha_{\rm s}$ is the power-law index of the high-density tail. The log-normal piece is found to well represent the isothermal density distribution of a low-density, turbulent ISM (e.g. Burkhart et al., 2009; Myers, 2015). The power-law tail at the high density, on the other hand, is expected to originate from compression by self-gravitating collapses and feedback (e.g. Slyz et al., 2005; Collins et al., 2011; Burkhart et al., 2017; Khullar et al., 2021; Appel et al., 2022; Buck et al., 2022; Appel et al., 2022; Kiihne et al., 2025). The continuity of $p(s)$ and its first derivative at $s=s_{\rm t}$ gives
@@
    $\alpha_{\rm s}$, the high-density slope, is found to be 11112222 by numerical simulations (e.g. Appel et al., 2022; Kiihne et al., 2025). We take $\alpha_{\rm s}=1.5$, which is consistent with the analytical expectation for the pressure-free (free-fall) collapse (Girichidis et al., 2014), and for the weak magnetic field and high compression we assumed in setting $f_{\rm s}$ (Collins et al., 2011).
@@ 3.2 The cooling and fragmentation of sub-clouds
Grey shades in Fig. 6 show the distributions of sub-cloud density at formation ($n_{\rm sc}$), for all sub-clouds formed in the main branches of central subhalos at $z=0$ with different halo masses. Depending on internal properties and environments, sub-clouds can have different fates: some will be destroyed by external feedback, while others will continue to contract, fragment and form stars at $n_{\rm sc}^{\prime}$. Only the most massive and/or densest sub-clouds can shield themselves from feedback and form sufficient amounts of stars to qualify as (the progenitors of) globular clusters. These will be discussed thoroughly in §§3.3, 3.4 and 3.5, and be synthesized into a coherent model in §4.1. Briefly, the aforementioned conditions, together with supersonic turbulence and metal-poor environment, give rise to two distinct formation channels that lead to two populations of GCs: Pop-I (the metal-rich population) and Pop-II (the metal-poor population).
@@ 3.4 Star formation in sub-clouds
This efficiency reaches unity at $\Sigma_{\rm sc}\approx 10^{3.2}\,{\rm M_{\odot}}{\rm pc}^{-2}$, comparable to the condition for sub-clouds with $\epsilon_{\rm w}\approx 1$. At high metallicity, the momentum injection from multiple scattering can effectively reduce the star formation. A more general form of Eq. (55) was originally proposed by Fall et al. (2010), and subsequently confirmed in hydrodynamical simulations and observations. In particular, Grudić et al. (2018) showed that the formula is accurate for a mixture of feedback processes, both momentum-driven and energy-driven, including radiation pressure, photo-ionization, stellar winds, etc. The formula is also consistent with various observations of proto-clusters/sub-clouds in the MW and the Large Magellanic Cloud (LMC; Mok et al., 2021).
@@ 4.2 Mass and age distributions
Figure 9: Mass functions of GCs. Different panels show the mass functions of GCs in central galaxies with different halo masses at $z=0$. Both in-situ and ex-situ GCs are included. In each panel, grey curve shows the birth-time mass function of all GCs formed in the history. Black curve shows the $z=0$ mass function of GCs survived after the mass loss due to dynamical evolution is incorporated, with dark shade showing the $1$-$\sigma$ range among individual galaxies and light shade showing the minimum to maximum. Red and blue curves show the Pop-I and Pop-II sub-components, respectively, of the black curve. See §4.1 for the details of modeling and sample selection. In the center panel, the histograms show the observed GC mass functions for MW (green) and M31 (purple), compiled by Rodriguez et al. (2023). Orange markers show the mass function of star clusters in M51, obtained from the star-cluster catalog of LEGUS (Messa et al., 2018). See §4.2 for a detailed discussion of this figure.
Figure 10: Age distributions of GCs. Different panels show the age distributions of GCs in central galaxies with different halo masses at $z=0$. Both in-situ and ex-situ GCs are included. In each panel, grey curve shows the distribution of the birth time of all GCs formed in the history. Black curve shows that of GCs survived until $z=0$ after the mass loss due to dynamical evolution is incorporated, with dark shade showing the $1$-$\sigma$ range among individual galaxies and light shade showing the minimum to maximum. Red and blue curves show the Pop-I and Pop-II sub-components, respectively, of the black curve. See §4.1 for the details of modeling and sample selection. In the center panel, the histograms show the results of the empirical model of Chen & Gnedin (2024) for MW (green) and M31 (purple). Their results for all and survived GCs are shown by dashed and solid histograms, respectively. Orange markers show the age distribution of star clusters more massive than $10^{5}{\rm M_{\odot}}$ in M51, obtained from the star-cluster catalog of LEGUS (Messa et al., 2018). See §4.2 for a detailed discussion of this figure.
@@
In the center panel of Fig. 9, we also include the observed mass functions of GCs in the Milky Way (green curve) and M31 (purple curve), compiled by Rodriguez et al. (2023, see their figure 3) using catalogs from Harris (1996, 2010) and Caldwell et al. (2011), respectively. The observed GC mass function of M31 has an amplitude about $0.5\,{\rm dex}$ higher than that of the MW, and appears more top-heavy. This may be explained partly by the more massive halo of M31, and partly by the fact that it has more recent mergers that can bring in ex-situ GCs as well as increase the in-situ formation of GCs (see the age distribution below; see also, e.g. §4.3 of Chen & Gnedin, 2024). The mass function of GCs predicted by our model (black curve) falls between those of MW and M31 over the entire range of $M_{\rm cls}$, indicating that the model is capable of reproducing the average GC mass function for MW-size galaxies. The observed mass functions of MW and M31 are well within the range of the model prediction (light shade), indicating that they are not outliers in the ensemble of MW-sized galaxies predicted by our model.
@@
The difference between MW and M31 in the mass and age distributions of GCs indicates the diversity of GC formation in galaxies with different states and assembly histories. To clearly see this, we consider another galaxy, M51 (NGC 5194), a spiral galaxy interacting with a massive companion, NGC 5195. The stellar mass of M51 is inferred to be $\approx 10^{10.6}{\rm M_{\odot}}$ (Leroy et al., 2008; Wei et al., 2021), slightly smaller than that of MW. We take the star-cluster candidates of the LEGUS survey for M51 (Messa et al., 2018), remove the contamination by selecting only those with machine-learning class 1 or 2, and exclude those with age less than $10\,{\rm Myr}$ to avoid unbound systems. The mass distribution of the star clusters so obtained is shown in the center panel of Fig. 9. We further restrict the sample to massive clusters with $M_{\rm cls}\geqslant 10^{5}{\rm M_{\odot}}$ so that they are age-complete (see Fig. 5 of Messa et al., 2018), and show the age distribution of these clusters in the center panel of Fig. 10. At $M_{\rm cls}\gtrsim 10^{5}{\rm M_{\odot}}$, the cluster mass function of M51 is much lower than those of MW, M31 and the model prediction. The age distribution shows that these massive clusters are mostly born at $z<0.5$ by recent star formation, and none of them are as old as those produced at $z\approx 5$, the peak of the GC-formation activity predicted by our model. Such a situation can be created if M51 forms very late, and is hosted by a halo with very stable dynamical environment in the early stage. The lack of galactic-scale turbulence thus suppresses the formation of old and massive GCs, and large amounts of GCs are produced only later by gas compression through, e.g. disk instabilities and/or external interactions (see the discussion in §3.1). Even considering the extreme ranges of the mass and age distributions predicted by our model (light shades in Figs. 9 and 10), the observed distributions of M51 are still outliers. This highlights the importance of explicitly modeling effects of galaxy interactions on the formation of GCs. Such interactations can complement and even dominate the source of turbulence in the ISM, and may produce a density distribution different from that based on the assumption of the single Mach number given by Eq. (36).